# Shuangrui Ding1  Rui Qian1  Xiaoyi Dong1,2  Pan Zhang2 Yuhang Zang2  Yuhang Cao2  Yuwei Guo1   Dahua Lin1   Jiaqi Wang2

### v2

@@ 4.3 Ablation Study
Table 6: Ablation on IoU threshold $\delta_{\text{iou}}$.
@@
0 80.0 76.6 83.4
0.3 80.8 77.1 84.5
0.7 80.2 76.6 83.8
0.9 77.8 74.3 81.3
Table 7: Ablation on uncertainty threshold $\delta_{\text{conf}}$.
[…]
0.5 80.4 76.7 83.7
2 80.8 77.1 84.5
5 80.5 76.9 84.1
Table 8: Ablation on modulation $[w_{\text{low}},w_{\text{high}}]$.
[…]
$\left[1,1\right]$ 80.2 76.5 83.8
$\left[0.95,1.05\right]$ 80.8 77.1 84.5
$\left[0.9,1.1\right]$ 80.5 76.9 84.1
4.4 Additional Results on VOT Benchmarks
To further evaluate the effectiveness of SAM2Long, we also apply it to three widely used single-object tracking benchmarks, namely LaSOT (Fan et al., 2019b), LaSOText (Fan et al., 2021), and GOT-10k (Huang et al., 2019). Without whistles and bells, we first generate an initial mask from the bounding box provided for the first frame using the SAM2 Mask decoder. Subsequently, all subsequent masks are transformed into bounding box format by identifying the minimum and maximum coordinates of their non-zero indices. For comparison, we include recent state-of-the-art methods specifically designed for VOT, including SAMURAI (Yang et al., 2024) and DAM2SAM (Videnovic et al., 2024).
Table 9: Performance comparison of SAM2Long with state-of-the-art VOT methods on LaSoT, LaSoText, and GOT-10k benchmarks. The results are reported in terms of AUC (Area Under the success rate Curve) for LaSoT and LaSoText, and AO (Average Overlap) for GOT-10k.
[…]
KeepTrack (Mayer et al., 2021) 67.1 48.2 -
TOMP (Mayer et al., 2022) 68.5 - -
DropTrack (Wu et al., 2023a) 71.8 52.7 75.9
SeqTrack (Chen et al., 2023) 72.5 50.7 74.8
MixFormer (Cui et al., 2022) 70.1 - 71.2
GRM-256 (Gao et al., 2023b) 69.9 - 73.4
ROMTrack (Cai et al., 2023) 71.4 51.3 74.2
OSTrack (Ye et al., 2022) 71.1 50.5 73.7
DiffusionTrack (Luo et al., 2024) 72.3 - 74.7
ODTrack (Zheng et al., 2024) 74.0 53.9 78.2
LORAT (Lin et al., 2025) 75.1 56.6 78.2
SAM2.1 (Ravi et al., 2024) 70.0 56.9 80.7
DAM2SAM (Videnovic et al., 2024) 75.1 60.9 81.1
SAMURAI (Yang et al., 2024) 74.2 61.0 81.7
SAM2Long (ours) 73.9 60.9 81.1
Based on the strong foundation of SAM2.1, we achieve competitive performance with state-of-the-art VOT methods. Specifically, our method, SAM2Long, delivers results that are comparable to the performance of SAMURAI and DAM2SAM, two recent methods specifically optimized for VOT.
4.5 Visualization
We present a qualitative comparison between SAM 2 and SAM2Long in Figure 3. SAM2Long demonstrates a significant reduction in segmentation errors, maintaining more accurate and consistent tracking of objects across various frames.
For example, in the second sequence of the second row, SAM 2 immediately loses track of the man of interest when occlusion happens. Although SAM2Long also loses track initially, its memory tree with multiple pathways enables it to successfully re-track the correct man later on. In another case, depicted in the third row where a group of people is dancing, SAM 2 initially tracks the correct person. However, when occlusion occurs, SAM 2 mistakenly switches to tracking a different, misleading individual. In contrast, SAM2Long handles this ambiguity effectively. Even during the occlusion, SAM2Long manages to resist the tracking error and correctly resumes tracking the original dancer when she reappears.
In conclusion, SAM2Long significantly improves SAM 2’s ability to handle object occlusion and reappearance, thereby enhancing its performance in long-term video segmentation.
5 Conclusion
In this paper, we introduce SAM2Long, a training-free enhancement to SAM 2 that alleviates its limitations in long-term video object segmentation. By employing a constrained tree memory structure with object-aware memory modulation, SAM2Long effectively mitigates error accumulation and improves robustness against occlusions, resulting in a more reliable segmentation process over extended periods. Extensive evaluations on six VOS benchmarks demonstrate that SAM2Long consistently outperforms SAM 2, especially in complex video scenarios. Notably, SAM2Long achieves up to a 5-point improvement in $\mathcal{J}\&\mathcal{F}$ scores on challenging long-term video benchmarks such SA-V and LVOS without requiring additional training or external parameters.
Although SAM2Long introduces significant improvements, there is still room for further enhancement. Future work could include fine-tuning the model on occlusion-heavy datasets using the memory architecture of SAM2Long. Additionally, exploring the semantic interactions between multiple objects within the same frame may offer valuable insights for more accurate segmentation, as the current method does not account for multi-object interactions in such scenarios.
@@ 5 Conclusion
Figure 3: Qualitative comparison between SAM 2 and SAM2Long, with GT (Ground Truth) provided for reference. A blue box is used to highlight incorrectly segmented objects, while a red box indicates missing objects. Best viewed when zoomed in.

### v3

@@ Appendix A Dataset Information
Table 10: Ablation study on uncertainty threshold $\delta_{\text{conf}}$.
$[w_{\text{low}},w_{\text{high}}]$ SAM 2 [0.6,1.4][0.6,1.4][ 0.6 , 1.4 ] [0.7,1.3][0.7,1.3][ 0.7 , 1.3 ] [0.8,1.2][0.8,1.2][ 0.8 , 1.2 ] [0.9,1.1][0.9,1.1][ 0.9 , 1.1 ] [0.95,1.05][0.95,1.05][ 0.95 , 1.05 ] [1,1][1,1][ 1 , 1 ]
SA-V 76.3 77.1 79.3 79.8 80.4 80.8 80.5
LVOS 83.0 77.0 78.8 82.9 85.0 85.4 84.9
Table 11: Ablation study on modulation weight $[w_{\text{low}},w_{\text{high}}]$.
SA-V val Single-object Multi-object Overall
# of seq 73 82 155
SAM 2.1 79.0 78.4 78.6
SAM2.1Long 80.6 (1.6% $\uparrow$) 81.3 (2.9% $\uparrow$) 81.1 (2.5% $\uparrow$)
Table 12: Performance comparison on single-object sequence and multi-object sequence. SAM2Long performs exceptionally well in both single-object and multi-object cases.
Appendix B More Ablation Study
Iou Threshold $\delta_{\text{iou}}$. The choice of the IoU threshold $\delta_{\text{iou}}$ is crucial for selecting frames with reliable object cues. As shown in Table 12, setting $0.1\geq\delta_{\text{iou}}\geq 0.7$ yields the competitive $&\mathcal{J}\&\mathcal{F}$, indicating an effective trade-off between filtering out poor-quality frames and retaining valuable segmentation information. In contrast, setting no quality requirement for masks ($\delta_{\text{iou}}=0$) lowers the score to 80.0, as unreliable frames with poor segmentation harm SAM 2. Conversely, an overly strict selection ($\delta_{\text{iou}}\geq 0.8$) further degrades performance by excluding important neighboring frames, forcing the model to rely on distant frames as memory.
Uncertainty Threshold $\delta_{\text{conf}}$. The uncertainty threshold $\delta_{\text{conf}}$ controls the selection of hypotheses under uncertain conditions. Our results in Table 12 indicate that setting $\delta_{\text{conf}}$ to 2 provides the highest $&\mathcal{J}\&\mathcal{F}$ score, indicating an optimal level for uncertainty handling. Lower values (e.g., 0) lead to suboptimal performance by committing to incorrect segmentations, causing error propagation. Higher values (e.g., 4) do not improve performance, indicating that beyond a certain threshold, the model efficiently relies on top-scoring masks without needing additional diversity.
@@ Appendix B More Ablation Study
Appendix C More Quantitive Comparion
In this section, we present a detailed comparison of SAM2.1 and SAM2Long across single-object and multi-object sequences, as shown in Table 12. Our experiments demonstrate that SAM2Long outperforms SAM2.1 in both single-object and multi-object scenarios. Specifically, SAM2Long achieves a 1.6% improvement in single-object sequences, a 2.9% improvement in multi-object sequences, and an overall 2.5% enhancement in performance. These results highlight SAM2Long’s robustness and effectiveness in various video segmentation tasks.
Appendix D More Visualization
We present additional comparisons between SAM2 and SAM2Long in Figure 4. SAM2Long significantly reduces segmentation errors, showing improved accuracy and consistency in object tracking across frames. Notably, in the Fast & Furious movie scene, SAM2Long successfully tracks the green car, even under challenging dynamic camera movements. Overall, SAM2Long offers substantial improvements over SAM2, especially in handling object occlusion and reappearance, leading to better performance in long-term video segmentation tasks.
@@ Appendix D More Visualization
Figure 4: Qualitative comparison between SAM 2 and SAM2Long, with GT (Ground Truth) provided for reference. The last row shows an in-wild case. Best viewed when zoomed in.